# Renormalization group approach to graphene bilayers

### v1

@@ I Introduction
Moreover, it has been shown that in the ordered, flat phase of two-dimensional polymerized membranes, which is controlled from the renormalization group (RG) point of view by an infrared fixed point [23, 24], one observes power-law behaviours for both the phonon–phonon and flexuron–flexuron correlation functions [23, 25, 19, 26]:
@@
On the numerical side, atomistic simulations have provided a complementary microscopic picture. Monte Carlo simulations of bilayer graphene by Zakharchenko et al. [28] have shown that the bending rigidity of a free bilayer is approximately twice that of a monolayer, and that there is a crossover from correlated to uncorrelated out-of-plane fluctuations of the two layers at a wave vector of order $q\sim 3\,\mathrm{nm}^{-1}$, highlighting the role of interlayer coupling in the long-wavelength flexural response.More recently, Herrero and Ramírez [49] have carried out extensive molecular-dynamics simulations of bilayer graphene over a wide temperature range, showing that the in-plane compression modulus of the bilayer is larger than that of a monolayer but decreases with increasing temperature, and that the onset of mechanical instability under in-plane compression is controlled by a size-dependent spinodal pressure and pronounced finite-size effects in the out-of-plane fluctuations and elastic constants. Their analysis of height fluctuations also indicates that out-of-plane corrugations are significantly reduced in the bilayer compared to a monolayer of the same lateral size and temperature, consistently with an enhanced effective bending rigidity. In addition, molecular-dynamics and Monte Carlo simulations of few-layer graphene and related van der Waals stacks reveal a rich thickness dependence of the flexural response, including a strong softening of the apparent bending stiffness due to interlayer shear and slip at large bending angles [50, 51]. Taken together, these studies indicate that the mechanical response of bilayer and multilayer membranes is governed by a nontrivial interplay between in-plane elasticity, interlayer coupling, and thermal fluctuations, and call for a theoretical framework capable of treating these ingredients on an equal footing and across scales.
On the theoretical side, only a few investigations have been carried out. A long-distance elastic action incorporating interlayer interactions was first proposed by de Andrés et al. in Ref. [52], and later derived more systematically by Mauri et al. in Ref. [45]. The latter authors studied this model within a SCSA framework and identified a crossover for the (bare) bending rigidity between two regimes: at short distances, a behaviour controlled by the in-plane elastic properties, and at long distances, a behaviour dominated by the bending rigidity of the two layers. This specific rigidity crossover comes in addition to the usual one already known for monolayers, namely the crossover between a quasi-harmonic regime at short distances and a strong-coupling, anharmonic regime at long distances, induced by nonlinear interactions between flexural modes.
@@ II.1.1 Action
In Eq. (3), Greek indices run from $1$ to $2$ and repeated indices are summed over. The action Eq. (3) contains four contributions: the first term describes the bending energy, with bending rigidity $\kappa$, while the second term represents an effective surface tension, with tension parameter $t$. The last two terms account for stretching and shear elastic energies arising from the fixed connectivity, and are characterized by the two Lamé coefficients $\lambda$ and $\mu$. Mechanical stability requires that $\kappa$, $\mu$, and the bulk modulus $B=\lambda+\mu$ are all positive. Note that in Eq. (3) the temperature has been absorbed into the definition of the field and of the coupling constants.
@@ II.2.1 Flat phase
We now turn to the field theoretical description of membrane bilayers composed of two monolayers 1 and 2 separated by a distance $\ell$. In doing so, we follow the constructions of Refs. [52, 45]. Each monolayer is described by a free action $S[{\bf R}_{i}]$, $i=1,2$, where $S[{\bf R}]$ is given by Eqs. (3), (5) or (6), and ${\bf R}_{i}$ ($i=1,2$) are the position fields of the two membranes, see Fig. 1.
@@
with $\bf n$} the normal to the monolayers
@@ II.2.2 Symmetries
Under a simultaneous translation $\mathbf{R}_{1,2}\to\mathbf{R}_{1,2}+\mathbf{C}$, with $\mathbf{C}$ an arbitrary three-dimensional vector, the center-of-mass coordinate $\mathbf{R}=\tfrac{1}{2}(\mathbf{R}_{1}+\mathbf{R}_{2})$ transforms as $\mathbf{R}\to\mathbf{R}+\mathbf{C}$, while the relative coordinate $\mathbf{S}=\mathbf{R}_{1}-\mathbf{R}_{2}$ remains invariant. Translational invariance therefore constrains the effective theory so that $\mathbf{R}$ can only appear through its derivatives, $\partial_{\alpha}\mathbf{R}$. The corresponding excitations are Goldstone modes associated with the spontaneous breaking of translational symmetry by the presence of the membrane, and are consequently massless. In contrast, because $\mathbf{S}$ is invariant under global translations, this symmetry does not restrict the form of local potentials $V(\mathbf{S})$ or couplings such as $({\bf S}^{2}-\ell^{2})\,u_{\alpha\alpha}$ that mix $\mathbf{S}$ with the in-plane strain of the center-of-mass field $\mathbf{R}$, see below. Fluctuations in $\mathbf{S}$ describe internal degrees of freedom, such as local variations of thickness or relative tilt between the two monolayers, and acquire a finite mass.
@@ II.2.3 Action
The explicit form of the interaction $S_{I}[{\bf R},{\bf S}]$ between the two membranes can be obtained from the above symmetry considerations, together with the requirement that, in order to preserve rotational invariance in the embedding space, it must be built only from scalar combinations of $\{\partial_{\alpha}{\bf R}\}$ and $\bf S$}. In addition, it should describe small fluctuations around the configuration Eq. (9). Accordingly, it can depend on the in-plane strain tensor $u_{\alpha\beta}$, the interlayer shear ${\bf S}\cdot\partial_{\alpha}{\bf R}$, and fluctuations of the interlayer distance, encoded in the term ${\bf S}^{2}-\ell^{2}$. The action compatible with all these symmetries is given by [45]
$\begin{split}S_{I}[{\bf R},{\bf S}]=\int\mathrm{d}^{2}x\,\bigg\{&\frac{g_{1}}{8\ell^{4}}\big({\bf S}^{2}-\ell^{2}\big)^{2}+\frac{g_{2}}{2\ell^{2}}\big({\bf S}\cdot\partial_{\alpha}{\bf R}\big)^{2}\\ &+\frac{g_{3}}{4\ell^{2}}\big({\bf S}^{2}-\ell^{2}\big)\,u_{\alpha\alpha}\bigg\}\,.\end{split}$ (10)
Note that here we have neglected anisotropic terms which are relevant for realistic graphene-based materials but are expected to be subdominant compared to the contributions retained in Eq. (10) (see Ref. [45]). The interaction action $S_{I}[{\bf R},{\bf S}]$ contains three terms. The first one is a restoring potential that drives ${\bf S}^{2}$ towards $\ell^{2}$. The second term is a tilt (director-tangent) coupling: it penalizes the projection of $\mathbf{S}$ onto the local tangent directions $\partial_{\alpha}\mathbf{R}$, and thus suppresses relative tilt of the interlayer director $\mathbf{S}$ with respect to the membrane’s tangent frame, stabilizing the alignment of $\mathbf{S}$ along the local normal. The last term is a volume-area coupling: it couples thickness fluctuations to areal strain, so that stretching ($u_{\alpha\alpha}>0$) favors thinning (${\bf S}^{2}<\ell^{2}$), while compression ($u_{\alpha\alpha}<0$) favors thickening (${\bf S}^{2}>\ell^{2}$), thereby encoding approximate local volume conservation.
@@
$S[$ $\boldsymbol{R},\boldsymbol{S}]=\int\mathrm{d}^{2}x\bigg\{\frac{\kappa}{2}\Big(2\,(\partial^{2}\boldsymbol{R})^{2}+\frac{1}{2}\,(\partial^{2}\boldsymbol{S})^{2}\Big)+\lambda\,u_{\alpha\alpha}^{2}$ (11)
$\hskip-11.38092pt+2\,\mu\,u_{\alpha\beta}^{2}+\frac{\lambda}{4}\,(\partial_{\alpha}\boldsymbol{R}.\partial_{\alpha}\boldsymbol{S})^{2}+\frac{\mu}{4}\,(\partial_{\alpha}\boldsymbol{S}.\partial_{\beta}\boldsymbol{R})^{2}$
$\hskip-11.38092pt+\frac{\lambda}{64}(\partial_{\alpha}\boldsymbol{S}.\partial_{\alpha}\boldsymbol{S})^{2}+\frac{\mu}{32}(\partial_{\alpha}\boldsymbol{S}.\partial_{\beta}\boldsymbol{S})^{2}+\frac{\lambda}{4}u_{\alpha\alpha}\,\partial_{\beta}\boldsymbol{S}.\partial_{\beta}\boldsymbol{S}$
$\hskip-11.38092pt+\frac{\mu}{2}u_{\alpha\beta}\,\partial_{\alpha}\boldsymbol{S}.\partial_{\beta}\boldsymbol{S}+\frac{\mu}{4}(\partial_{\alpha}\boldsymbol{R}.\partial_{\beta}\boldsymbol{S})(\partial_{\alpha}\boldsymbol{S}.\partial_{\beta}\boldsymbol{R})$
$\hskip-11.38092pt+\frac{g_{1}}{8\ell^{4}}(\boldsymbol{S}^{2}-\ell^{2})^{2}+\frac{g_{2}}{2\ell^{2}}(\boldsymbol{S}.\partial_{\alpha}\boldsymbol{R})^{2}+\frac{g_{3}}{4\ell^{2}}\big({\bf S}^{2}-\ell^{2})u_{\alpha\alpha}\bigg\}.$
@@
Following and drawing on the works of de Andrés et al. [52] and Mauri et al. [45], we will focus, among the three interaction terms, on the one responsible for the rigidity crossover, namely the coupling proportional to $g_{2}$ that we take very large. Moreover we will take the limit $g_{1}\to\infty$, which fixes the interlayer separation to $|{\bf S}|=\ell$. In this limit, the term proportional to $g_{3}$ does not contribute, since it is multiplied by ${\bf S}^{2}-\ell^{2}$.
@@ III.1 Methodology
$\Delta S_{k}[{\bf R}]=\frac{1}{2}\int_{\bf q}R_{k,ij}({\bf q})\,R_{i}({\bf q})\,R_{j}(-{\bf q})\,$, (13)
where $R_{k,ij}({\bf q})$ ($i,j=1,\dots,3$) is a $\bf q$}-dependent cut-off function, which is very often taken to be diagonal, $R_{k,ij}({\bf q})=R_{k}({\bf q})\delta_{ij}$. The partition function in the presence of a source ${\bf J}({\bf q})$ then reads
$Z_{k}[{\bf J}]=\int\mathcal{D}{\bf R}\,\exp\bigg[$ $-S[{\bf R}]-\Delta S_{k}[{\bf R}]$ (14)
$+\int_{\bf q}{\bf J}({\bf q})\cdot{\bf R}(-{\bf q})\bigg],$
with $\int_{\bf q}\equiv\int\mathrm{d}^{2}{\bf q}/(2\pi)^{2}$. The role of the term Eq. (13) is to suppress low-wavevector modes so that only high-wavevector modes are effectively integrated out in Eq. (14). In order to implement this program and ensure that $\Gamma_{k}$ satisfies the limits Eq. (12), the cut-off function $R_{k}({\bf q})$ must satisfy the following properties:
@@
  • $R_{k}({\bf q})\sim k^{\alpha}$ with $\alpha>0$ when $|{\bf q}|\ll k$: low-wavevector modes acquire an effective “mass” of order $k^{\alpha}$ and thus effectively decouple.
  • $R_{k}({\bf q})\to 0$ when $|{\bf q}|\gg k$: high-wavevector modes are essentially unaffected by the regulator term.
@@
  • $R_{k}({\bf q})\to 0$ when $k\to 0$: the regulator term vanishes when all fluctuations have been integrated out, so that the mass term no longer plays any role.
  • $R_{k}({\bf q})\to\Lambda^{\alpha}$ when $k\to\Lambda$ (with $\Lambda$ generally taken infinite): at the lattice scale, $R_{k}({\bf q})$ acts as a very large mass for all modes. They are effectively “frozen” and do not contribute to the functional integral Eq. (14).
@@
$R_{k}({\bf q})=Z_{k}\big(k^{\alpha}-|{\bf q}|^{\alpha}\big)\,\Theta\big(k^{\alpha}-|{\bf q}|^{\alpha}\big),$ (16)
@@
Figure 2: A typical shape of the cut-off function $R_{k}({\bf q})$.
@@
with $W_{k}[{\bf J}]=\log Z_{k}[{\bf J}]$, where we have set $k_{B}T=1$, and where $\bf r$} is the expectation value of the microscopic field $\bf R$} in the presence of the source $\bf J$}:
@@
With this definition, together with the properties of the cut-off function $R_{k}({\bf q})$ specified above, it is straightforward to check that $\Gamma_{k}$ satisfies the limits Eq. (12).
@@
$\partial_{t}\Gamma_{k}[{\bf r}]$ $=\frac{1}{2}\int_{\bf q}\partial_{t}R_{k,ij}({\bf q})\left(\Gamma_{k,ji}^{(2)}[{\bf r};{\bf q},{-{\bf q}}]+R_{k,ji}({\bf q})\right)^{-1}$ (19)
$\equiv\frac{1}{2}{\hbox{Tr}}\left[\partial_{t}R_{k}\left(\Gamma_{k}^{(2)}+R_{k}\right)^{-1}\right]$
@@
$\partial_{t}\Gamma_{k}[{\bf r}]=\frac{1}{2}\widehat{\partial}_{t}{\hbox{Tr}}\ln\left[\Gamma_{k}^{(2)}+R_{k}\right]$ (20)
with $\widehat{\partial}_{t}\equiv\partial_{t}R_{k}({\bf q}){\partial_{R_{k}({\bf q})}}$.
@@
Eqs. (19) and (20) govern the evolution of $\Gamma_{k}$ as $k$ is lowered, i.e. as fluctuations with lower and lower wavevectors are successively included. In these equations, the cut-off function $R_{k}$ appears both in the numerator of the integrand, through $\partial_{t}R_{k}$, and in the denominator as a “mass term” in the propagator. In the former case, combined with the generic shape of the regulated propagator (see Fig. 2), this ensures that only modes with wavevectors in a shell around $k$ contribute effectively to the RG flow, in accordance with the Wilsonian picture. In particular, $R_{k}$ controls the ultraviolet (UV) sector. In the latter case, $R_{k}$ acts as an infrared (IR) regulator, which guarantees that even at criticality any singular behaviour is shifted to vanishing wavevector in the limit $k\to 0$. This mechanism works provided the exponent $\alpha$ in $k^{\alpha}$ is chosen appropriately: for scalar theories with $O(N)$ symmetry and a standard kinetic term of order $\partial^{2}$, one takes $\alpha=2$, whereas for crystalline membranes, whose kinetic term is of order $\partial^{4}$, one chooses $\alpha=4$ [40]. With this choice, one can safely investigate any phase or phase transition by taking the limit $k\to 0$.

### v2

# Renormalization group approach to the elastic properties of graphene bilayers

@@ I Introduction
Moreover, it has been shown that in the ordered, flat phase of two-dimensional polymerized membranes, which is controlled from the renormalization group (RG) point of view by an infrared fixed point [23, 24], one observes power-law behaviors for both the phonon–phonon and flexuron–flexuron correlation functions [23, 25, 19, 26]:
@@
On the numerical side, atomistic simulations have provided a complementary microscopic picture. Monte Carlo simulations of bilayer graphene by Zakharchenko et al. [28] have shown that the bending rigidity of a free bilayer is approximately twice that of a monolayer, and that there is a crossover from correlated to uncorrelated out-of-plane fluctuations of the two layers at a wave vector of order $q\sim 3\,\mathrm{nm}^{-1}$, highlighting the role of interlayer coupling in the long-wavelength flexural response.More recently, Herrero and Ramírez [49] have carried out extensive molecular-dynamics simulations of bilayer graphene over a wide temperature range, showing that the in-plane compression modulus of the bilayer is larger than that of a monolayer but decreases with increasing temperature, and that the onset of mechanical instability under in-plane compression is controlled by a size-dependent spinodal pressure and pronounced finite-size effects in the out-of-plane fluctuations and elastic constants. Their analysis of height fluctuations also indicates that out-of-plane corrugations are significantly reduced in the bilayer compared to a monolayer of the same lateral size and temperature, consistently with an enhanced effective bending rigidity. In addition, molecular-dynamics and Monte Carlo simulations of few-layer graphene and related van der Waals stacks reveal a rich thickness dependence of the flexural response, including a strong softening of the apparent bending stiffness because of interlayer shear and slip at large bending angles [50, 51]. Taken together, these studies indicate that the mechanical response of bilayer and multilayer membranes is governed by a nontrivial interplay between in-plane elasticity, interlayer coupling, and thermal fluctuations, and call for a theoretical framework capable of treating these ingredients on an equal footing and across scales.
On the theoretical side, only a few investigations have been carried out. A long-distance elastic action incorporating interlayer interactions was first proposed by de Andrés et al. in Ref. [52], and later derived more systematically by Mauri et al. in Ref. [45]. The latter authors studied this model within a SCSA framework and identified a crossover for the (bare) bending rigidity between two regimes: at short distances, a behavior controlled by the in-plane elastic properties, and at long distances, a behavior dominated by the bending rigidity of the two layers. This specific rigidity crossover comes in addition to the usual one already known for monolayers, namely the crossover between a quasi-harmonic regime at short distances and a strong-coupling, anharmonic regime at long distances, induced by nonlinear interactions between flexural modes.
@@ II.1.1 Action
In Eq. (3), Greek indices run from $1$ to $2$ and repeated indices are summed over. More generally, Einstein’s summation convention applies throughout, unless stated otherwise. The action Eq. (3) contains four contributions: the first term describes the bending energy, with bending rigidity $\kappa$, while the second term represents an effective surface tension, with tension parameter $t$. The last two terms account for stretching and shear elastic energies arising from the fixed connectivity, and are characterized by the two Lamé coefficients $\lambda$ and $\mu$. Mechanical stability requires that $\kappa$, $\mu$, and the bulk modulus $B=\lambda+\mu$ are all positive. Note that in Eq. (3) the temperature has been absorbed into the definition of the field and of the coupling constants.
@@ II.2.1 Flat phase
We now turn to the field theoretical description of membrane bilayers composed of two monolayers 1 and 2 separated by a distance $\ell$. In doing so, we follow the constructions of Refs. [52, 45]. Each monolayer is described by a free action $S[{\bf R}_{i}]$, $i=1,2$, where $S[{\bf R}]$ is given by Eqs. (3), (5) or (6), and ${\bf R}_{i}$, $i=1,2$, are the position fields of the two membranes, see Fig. 1.
@@
with $\bf n$} the normal to the monolayers.
@@ II.2.2 Symmetries
Under a simultaneous translation $\mathbf{R}_{1,2}\to\mathbf{R}_{1,2}+\mathbf{C}$, with $\mathbf{C}$ an arbitrary three-dimensional vector, the center-of-mass coordinate $\mathbf{R}=\tfrac{1}{2}(\mathbf{R}_{1}+\mathbf{R}_{2})$ transforms as $\mathbf{R}\to\mathbf{R}+\mathbf{C}$, while the relative coordinate $\mathbf{S}=\mathbf{R}_{1}-\mathbf{R}_{2}$ remains invariant. Translational invariance therefore constrains the effective theory so that $\mathbf{R}$ can only appear through its derivatives, $\partial_{\alpha}\mathbf{R}$. The corresponding excitations are Goldstone modes associated with the spontaneous breaking of translational symmetry by the presence of the membrane, and are consequently massless. In contrast, because $\mathbf{S}$ is invariant under global translations, this symmetry does not restrict the form of local potentials $V(\mathbf{S})$ or couplings such as $({\bf S}^{2}-\ell^{2})\,u_{\alpha\alpha}$ that mix $\mathbf{S}$ with the center-of-mass field $\mathbf{R}$ strain tensor $u_{\alpha\beta}$, see below. Fluctuations in $\mathbf{S}$ describe internal degrees of freedom, such as local variations of thickness or relative tilt between the two monolayers, and acquire a finite mass.
@@ II.2.3 Action
The explicit form of the interaction $S_{I}[{\bf R},{\bf S}]$ between the two membranes can be obtained from the above symmetry considerations, together with the requirement that, in order to preserve rotational invariance in the embedding space, it must be built only from scalar combinations of $\{\partial_{\alpha}{\bf R}\}$ and $\bf S$}. In addition, it should describe small fluctuations around the configuration Eq. (9). Accordingly, it can depend on the center-of-mass field $\mathbf{R}$ strain tensor $u_{\alpha\beta}$, the interlayer shear ${\bf S}\cdot\partial_{\alpha}{\bf R}$, and fluctuations of the interlayer distance, encoded in the term ${\bf S}^{2}-\ell^{2}$. The action compatible with all these symmetries is given by [45]
$\begin{split}S_{I}[{\bf R},{\bf S}]=\int\mathrm{d}^{2}x\,\bigg\{&\frac{g_{1}}{8\ell^{4}}\big({\bf S}^{2}-\ell^{2}\big)^{2}+\frac{g_{2}}{2\ell^{2}}\big({\bf S}\cdot\partial_{\alpha}{\bf R}\big)^{2}\\ &+\frac{g_{3}}{4\ell^{2}}\big({\bf S}^{2}-\ell^{2}\big)\,u_{\alpha\alpha}\bigg\}\end{split}$ (10)
with $u_{\alpha\beta}=1/2(\partial_{\alpha}{\bf R}.\partial_{\beta}{\bf R}-\partial_{\alpha}{\bf R}_{0}.\partial_{\beta}{\bf R}_{0})$. Note that here we have neglected anisotropic terms which are relevant for realistic graphene-based materials but are expected to be subdominant compared to the contributions retained in Eq. (10) (see Ref. [45]). The interaction action $S_{I}[{\bf R},{\bf S}]$ contains three terms. The first one is a restoring potential that drives ${\bf S}^{2}$ towards $\ell^{2}$. The second term is a tilt (director-tangent) coupling: it penalizes the projection of $\mathbf{S}$ onto the local tangent directions $\partial_{\alpha}\mathbf{R}$, and thus suppresses relative tilt of the interlayer director $\mathbf{S}$ with respect to the membrane’s tangent frame, stabilizing the alignment of $\mathbf{S}$ along the local normal. The last term is a volume-area coupling: it couples thickness fluctuations to areal strain, so that stretching ($u_{\alpha\alpha}>0$) favors thinning (${\bf S}^{2}<\ell^{2}$), while compression ($u_{\alpha\alpha}<0$) favors thickening (${\bf S}^{2}>\ell^{2}$), thereby encoding approximate local volume conservation.
@@
$S[$ ${\bf R},{\bf S}]=\int\mathrm{d}^{2}x\bigg\{\frac{\kappa}{2}\Big(2\,(\partial^{2}{\bf R})^{2}+\frac{1}{2}\,(\partial^{2}{\bf S})^{2}\Big)+\lambda\,u_{\alpha\alpha}^{2}$ (11)
$\hskip-11.38092pt+2\,\mu\,u_{\alpha\beta}^{2}+\frac{\lambda}{4}\,(\partial_{\alpha}{\bf R}.\partial_{\alpha}{\bf S})^{2}+\frac{\mu}{4}\,(\partial_{\alpha}{\bf S}.\partial_{\beta}{\bf R})^{2}$
$\hskip-11.38092pt+\frac{\lambda}{64}(\partial_{\alpha}{\bf S}.\partial_{\alpha}{\bf S})^{2}+\frac{\mu}{32}(\partial_{\alpha}{\bf S}.\partial_{\beta}{\bf S})^{2}+\frac{\lambda}{4}u_{\alpha\alpha}\,\partial_{\beta}{\bf S}.\partial_{\beta}{\bf S}$
$\hskip-11.38092pt+\frac{\mu}{2}u_{\alpha\beta}\,\partial_{\alpha}{\bf S}.\partial_{\beta}{\bf S}+\frac{\mu}{4}(\partial_{\alpha}{\bf R}.\partial_{\beta}{\bf S})(\partial_{\alpha}{\bf S}.\partial_{\beta}{\bf R})$
$\hskip-11.38092pt+\frac{g_{1}}{8\ell^{4}}({\bf S}^{2}-\ell^{2})^{2}+\frac{g_{2}}{2\ell^{2}}({\bf S}.\partial_{\alpha}{\bf R})^{2}+\frac{g_{3}}{4\ell^{2}}\big({\bf S}^{2}-\ell^{2})u_{\alpha\alpha}\bigg\}.$
@@
Following and drawing on the works of de Andrés et al. [52] and Mauri et al. [45], we will focus, among the three interaction terms, on the one responsible for the rigidity crossover, namely the coupling proportional to $g_{2}$. Moreover we will take the limit $g_{1}\to\infty$, which fixes the interlayer separation to $|{\bf S}|=\ell$. In this limit, the term proportional to $g_{3}$ does not contribute, since it is multiplied by ${\bf S}^{2}-\ell^{2}$.
@@ III Nonperturbative renormalization group approach to polymerized membranes
For readers already familiar with the NPRG formalism and with its application to monolayer membranes, Secs. III A and III B may be skipped on a first reading, and one can proceed directly to Sec. III C where the bilayer flow equations are derived.
@@ III.1 Methodology
$S_{k}[{\bf R}]=\frac{1}{2}\int_{\bf q}{\cal R}_{k,ij}({\bf q})\,R_{i}({\bf q})\,R_{j}(-{\bf q})\,$, (13)
with $\int_{\bf q}\equiv\int\mathrm{d}^{2}q/(2\pi)^{2}$. In Eq.(13), ${\cal R}_{k,ij}({\bf q})$ ($i,j=1,\dots,3$) is a $\bf q$}-dependent cut-off function, which is very often taken to be diagonal, ${\cal R}_{k,ij}({\bf q})={\cal R}_{k}({\bf q})\delta_{ij}$. The partition function in the presence of a source ${\bf J}({\bf q})$ then reads
${\cal Z}_{k}[{\bf J}]=\int\mathcal{D}{\bf R}\,\exp\bigg[$ $-S[{\bf R}]-\Delta S_{k}[{\bf R}]$ (14)
$+\int_{\bf q}{\bf J}({\bf q})\cdot{\bf R}(-{\bf q})\bigg]\,.$
The role of the term Eq. (13) is to suppress low-wavevector modes so that only high-wavevector modes are effectively integrated out in Eq. (14). In order to implement this program and ensure that $\Gamma_{k}$ satisfies the limits Eq. (12), the cut-off function ${\cal R}_{k}({\bf q})$ must satisfy the following properties:
@@
  • ${\cal R}_{k}({\bf q})\sim k^{\rho}$ with $\rho>0$ when $|{\bf q}|\ll k$: low-wavevector modes acquire an effective “mass” of order $k^{\rho}$ and thus effectively decouple.
  • ${\cal R}_{k}({\bf q})\to 0$ when $|{\bf q}|\gg k$: high-wavevector modes are essentially unaffected by the regulator term.
@@
  • ${\cal R}_{k}({\bf q})\to 0$ when $k\to 0$: the regulator term vanishes when all fluctuations have been integrated out, so that the mass term no longer plays any role.
  • ${\cal R}_{k}({\bf q})\to\Lambda^{\rho}$ when $k\to\Lambda$ (with $\Lambda$ generally taken infinite): at the lattice scale, ${\cal R}_{k}({\bf q})$ acts as a very large mass for all modes. They are effectively “frozen” and do not contribute to the functional integral Eq. (14).
@@
${\cal R}_{k}({\bf q})=Z_{k}\big(k^{\rho}-|{\bf q}|^{\rho}\big)\,\Theta\big(k^{\rho}-|{\bf q}|^{\rho}\big),$ (16)
@@
Figure 2: A typical shape of the cut-off function ${\cal R}_{k}({\bf q})$.
@@
with $W_{k}[{\bf J}]=\log{\cal Z}_{k}[{\bf J}]$, where we have set $k_{B}T=1$, and where $\bf r$} is the expectation value of the microscopic field $\bf R$} in the presence of the source $\bf J$}:
@@
With this definition, together with the properties of the cut-off function ${\cal R}_{k}({\bf q})$ specified above, it is straightforward to check that $\Gamma_{k}$ satisfies the limits Eq. (12).
@@
$\partial_{t}\Gamma_{k}[{\bf r}]$ $=\frac{1}{2}\int_{\bf q}\partial_{t}{\cal R}_{k,ij}({\bf q})\left(\Gamma_{k,ji}^{(2)}[{\bf r};{\bf q},{-{\bf q}}]+{\cal R}_{k,ji}({\bf q})\right)^{-1}$ (19)
$\equiv\frac{1}{2}{\hbox{Tr}}\left[\partial_{t}{\cal R}_{k}\left(\Gamma_{k}^{(2)}+{\cal R}_{k}\right)^{-1}\right]$
@@
$\partial_{t}\Gamma_{k}[{\bf r}]=\frac{1}{2}\widehat{\partial}_{t}{\hbox{Tr}}\ln\left[\Gamma_{k}^{(2)}+{\cal R}_{k}\right]$ (20)
with $\widehat{\partial}_{t}\equiv\partial_{t}{\cal R}_{k}({\bf q}){\partial_{{\cal R}_{k}({\bf q})}}$.
@@
Eqs. (19) and (20) govern the evolution of $\Gamma_{k}$ as $k$ is lowered, i.e. as fluctuations with lower and lower wavevectors are successively included. In these equations, the cut-off function ${\cal R}_{k}$ appears both in the numerator of the integrand, through $\partial_{t}{\cal R}_{k}$, and in the denominator as a “mass term” in the propagator. In the former case, combined with the generic shape of the regulated propagator (see Fig. 2), this ensures that only modes with wavevectors in a shell around $k$ contribute effectively to the RG flow, in accordance with the Wilsonian picture. In particular, ${\cal R}_{k}$ controls the ultraviolet (UV) sector. In the latter case, ${\cal R}_{k}$ acts as an infrared (IR) regulator, which guarantees that even at criticality any singular behavior is shifted to vanishing wavevector in the limit $k\to 0$. This mechanism works provided the exponent $\rho$ in $k^{\rho}$ is chosen appropriately: for scalar theories with $O(N)$ symmetry and a standard kinetic term of order $\partial^{2}$, one takes $\rho=2$, whereas for crystalline membranes, whose kinetic term is of order $\partial^{4}$, one chooses $\rho=4$ [40]. With this choice, one can safely investigate any phase or phase transition by taking the limit $k\to 0$.